Grade 11Physical Chemistry

The wavelength of a microscopic particle of mass 9.1*10'-34 is 182nm it's kinetic energy in J is

S Joycee
6 Years agoGrade 11

1 Answer

Vikas TU
6 Years ago
Dear student 
Wavelength = Lambda = h/mv 
182 * 10^-9 = (6.63 * 10 ^-34)/9.1 *10^-31 * v 
Calculate v from here 
KE = 1/2 mv^2 
KE will be 7.28×10^−24 J